archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    The SFR collect customers will switch onto the standby LNS whereas 15 minutes downtime is expected for FT collect Date Friday 20 December 2013 07 21AM Reason for closing Done Comments 14 Comment by OVH Wednesday 18 December 2013 00 17AM upgrading bootrom Comment by OVH Wednesday 18 December 2013 00 17AM 1st reboot Comment by OVH Wednesday 18 December 2013 00 20AM bootrom updated waiting for act stdy redundancy Comment by OVH Wednesday 18 December 2013 01 27AM now act standby supervision cards are in sync now downloading the new release into alternate slot Comment by OVH Wednesday 18 December 2013 01 27AM Erasing the alternate release Comment by OVH Wednesday 18 December 2013 01 27AM Installing the new release Comment by OVH Wednesday 18 December 2013 01 27AM Installation completed successfully now upgrading release Comment by OVH Wednesday 18 December 2013 01 27AM 2nd and last reboot Comment by OVH Wednesday 18 December 2013 01 28AM LNS up to date verifying setting Comment by OVH Wednesday 18 December 2013 01 28AM checking fpga version reload card 10 fpga Comment by OVH Wednesday 18 December 2013 01 28AM all fpgas are up to date Comment by OVH Friday 20

    Original URL path: http://status.ovh.es/?do=details&id=5989 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    Core Status CLOSED Percent Complete Details We re going to update the configuration of the Paris and Roubaix LNS tonight This will impact on the functioning of IP blocks 29 during the intervention Date Tuesday 17 December 2013 12 47PM Reason for closing Done Comments 3 Comment by OVH Thursday 12 December 2013 09 21AM we will start with the RBX LNS Comment by OVH Thursday 12 December 2013 09

    Original URL path: http://status.ovh.es/?do=details&id=5962 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    issue on the FT collect port in Roubaix We are contacting Orange Date Thursday 12 December 2013 15 38PM Reason for closing Done Comments 2 Comment by OVH Thursday 12 December 2013 13 30PM The national FT collect and the

    Original URL path: http://status.ovh.es/?do=details&id=5968 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Category Backend Core Status CLOSED Percent Complete Details An intervention on the fibre optic tubes is reaching the NRA The NRA will not be accessible while the works are being carried out The rest of the loop is not affected

    Original URL path: http://status.ovh.es/?do=details&id=5963 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    dept 31 Attached to Project ADSL SDSL Task Type Incident Category Backend Core Status CLOSED Percent Complete Details There is an incident on the SFR devices in département 31 A ticket has been opened Date Thursday 12 December 2013 12

    Original URL path: http://status.ovh.es/?do=details&id=5961 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    on Tuesday December 10th on the Lyon optical loop in order to insert chromatic dispersion compensation modules Securisation will be affected during insertion of the module and cables No service interruption is expected Date Tuesday 10 December 2013 17 05PM Reason for closing Done Comments 3 Comment by OVH Tuesday 10 December 2013 10 44AM We re starting the intervention at the PAR69 site Comment by OVH Tuesday 10 December

    Original URL path: http://status.ovh.es/?do=details&id=5949 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    for closing Done Comments 3 Comment by OVH Monday 09 December 2013 11 38AM There is no optical continuity on site A technician is on route to site Comment by OVH Monday 09 December 2013 12 29PM The site is operational again the fibre operateur confirmed an incident on the optic network between 11 00 and 11 15 with repair at 11 45 We are keeping the ticket open as

    Original URL path: http://status.ovh.es/?do=details&id=5942 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Category Backend Core Status CLOSED Percent Complete Details This is an issue on the collect in Départements 21 91 and 83 A ticket has been opened Date Wednesday 04 December 2013 14 39PM Reason for closing Done Comments 1 Comment

    Original URL path: http://status.ovh.es/?do=details&id=5907 (2015-10-15)
    Open archived version from archive



  •